### v1

Abstract
The European legislature has proposed the Digital Services Act (DSA) and Artificial Intelligence Act (AIA) to regulate platforms and Artificial Intelligence (AI) products. We review to what extent third-party audits are part of both laws and to what extent access to models and data is provided. By considering the value of third-party audits and third-party data access in an audit ecosystem, we identify a regulatory gap in that the Artificial Intelligence Act does not provide access to data for researchers and civil society. Our contributions to the literature include: (1) Defining an AI audit ecosystem that incorporates compliance and oversight. (2) Highlighting a regulatory gap within the DSA and AIA regulatory framework, preventing the establishment of an AI audit ecosystem. (3) Emphasizing that third-party audits by research and civil society must be part of that ecosystem and demand that the AIA include data and model access for certain AI products. We call for the DSA to provide NGOs and investigative journalists with data access to platforms by delegated acts and for adaptions and amendments of the AIA to provide third-party audits and data and model access at least for high-risk systems to close the regulatory gap. Regulations modeled after European Union AI regulations should enable data access and third-party audits, fostering an AI audit ecosystem that promotes compliance and oversight mechanisms.
Keywords:
Auditing, Accountability, Data Access, AI Act, DSA, Regulation
1 Introduction
Artificial Intelligence (AI) technologies can cause serious harms to individuals, groups, and society. These technologies and products include Machine Learning (ML)-based algorithmic decision-making (ADM) systems, social media recommendation systems and general-purpose AI including foundational and generative AI. For instance, discriminatory practices and opaqueness in the welfare fraud ADM of the city of Rotterdam were uncovered [17]. It was demonstrated that YouTube [74] progresses towards extreme right-winged content on the platform. Furthermore, we have seen harmful biases in ML models such as stable diffusion [59], face recognition [13] or ChatGPT [72]. The European legislature has responded with different regulatory approaches to address these risks. For example, it has proposed regulating very large online platforms [^1] Not restricted to AI-driven algorithms (VLOPs) and very large online search engines (VLOSEs) via the Digital Services Act (DSA), as well as AI-driven products, services, and systems in the territory of the European Union (EU) via the Artificial Intelligence Act (AIA).
Among other regulatory tools, internal or external systematic evaluations of an algorithm can create public accountability by uncovering problematic machine behavior [70, 8, 68] and are to different extends part of the DSA and AIA. It has been argued that the EU regulation thereby de facto sketches an EU-wide ecosystem for auditing AI systems (see e.g. [56]). However, others have argued that, as enforcement mechanisms for “high-risk” AI systems, self-assessments before deployment (first- and second-party ex ante audits) and post-market monitoring (first- and second-party ex post audits) do not provide sufficient oversight [21]. Therefore, some have raised the question of the extent to which audits and impact evaluations by third parties (regulator, research, and civil society) are necessary for sufficient oversight [21, 44]. As this paper will show, access to the model and training data of the algorithmic system is crucial to facilitate third-party auditing. In this respect, AlgorithmWatch’s concern that this third party access is lacking in the AIA[4] should be taken seriously. The DSA includes third-party access by vetted researchers. However, others have argued that companies could leverage their market power against third-party auditors (i.e. audit capture) [46].
In general, researchers wonder “who will audit internal auditors?” [19, 3, 21] Therefore, we analyze whether the DSA and the AIA can create an EU audit ecosystem. To do so, we examine the importance of third-party audits by researchers and civil society. We find that third-party audits are essential for the establishment of an AI auditing ecosystem in the EU that ensures accountability by compliance and oversight. We argue that the inclusion of such audits must be strengthened by considering the existing provisions of the DSA and the proposed provisions of the AIA. The question of how regulatory inspection should be included unfortunately goes beyond the scope of the paper, although researchers have justifiably insisted on the importance of this issue (see e.g. [36]).
We contribute to the existing literature concerning third-party audits in three ways: (1) We define an AI audit ecosystem that accounts for compliance and oversight. (2) We demonstrate the existence of a regulatory gap as the DSA and AIA will not lead to an AI audit ecosystem. (3) We emphasize that third-party audits by research and civil society have to be part of that ecosystem and we demand that the AIA should include data and model access for certain AI products.
First, we introduce the terms related to algorithmic audits and the different types of audits (section 2). We then analyze the current regulatory framework on third-party audits (section 3). The following section addresses third-party audit case studies and defines an AI audit ecosystem that ensures compliance and oversight. Furthermore, we show that third-party audits by researchers and civil society are vital for an AI audit ecosystem (section 4). Finally, we show that the current EU legislation – the DSA and AIA in combination – will not be sufficient to create an AI audit ecosystem (section 5).
2 Algorithm Audits: Terminology and Key Properties
In this section, our primary objective is to establish a robust foundation of terminology and key attributes inherent to audits. Consequently, we will distinguish audit types, examine the scope of audit types, and underline the role of sociotechnical system thinking in audits. These examinations are conducted in preparation for the analysis of the role of third parties in the DSA and AIA. Additionally, we will revisit these audit types in the discussion of creating a diverse AI audit ecosystem.
The term “audit”  [^2] Sandvig et al. [76] have noted that ”although the word ’audit’ may evoke financial accounting, the original audit studies were developed by government economists to detect racial discrimination in housing.” Later, they were introduced in the social science as field experiments to test for discriminatory behavior [29] is used in various ways. For algorithmic audits [^3] Audit and algorithm audit will be used interchangeably in the following, some (e.g. Koshiyama et al. [44]) have provided a more technical definition of an algorithm audit, which is related to verification and compliance, and others (e.g. Sandvig et al. [76] and Metaxa et al. [53]) have used the term to describe a narrowly but still systematically targeted test of a particular aspect of a system (e.g., bias), originating from social science “audit studies.”
Two examples demonstrate the various ways that the term has been used. The first, Bandy [8] identifies an audit research gap regarding Twitter audits. Twitter is the most researched social media platform due to its formerly openly-accessible API [62]. This suggests a conceptual problem: What distinguishes research on Twitter and an from an audit conducted on Twitter? In the second example Metaxa et al. [53] have defined audits that occur through systematic external evaluations by third parties as a means of activism with internal knowledge of the process or system being studied.
We, however, will use audits as an umbrella term for empirical studies that investigate algorithmic systems and AI systems in particular and evaluate performance relative to expected behavior, as part of a broader accountability process [70]. Bandy [8] captures this broad definition, defining an algorithmic audit as “empirical study investigating a public algorithmic system for potential problematic behaviour” that includes the mentioned notions and is agnostic to the time of the audit (before or after the deployment). “Potential problematic behavior” refers to behavior with the potential to cause harm and includes the potential harms described by a sociotechnical AI system. ”Empirical studies” refers to qualitative and quantitative, as system expectations can be articulated and assessed in either mode [70].
Bandy [8], who performed a systematic review of algorithm audits, notes that racial discrimination remains central to algorithm audits. Although so-called bias audits can account for algorithmic discrimination based on other protected attributes, such as age, gender, socioeconomic status, religion, and intersectional identities. In addition to such bias audits that can account for representational and allocative harms, other harm potentials are possible from AI systems. Therefore, audits might have different scopes to account for these harms, which include quality of service harms, interpersonal harms, and social system harms (e.g. Shelby et al. [79]). To account for these harms, audits like robustness audits, interpretability and explainability audits, security audits and privacy audits and can be conducted [44].
Based on Kak and West [39], the following definitions are used in this paper:
  • first-party audits: internal audits which are conducted in an organization that developed the AI system based on specific metrics, toolkits, and requirements.
  • second-party audits: conducted by contractors with the developer (audits-as-a-service), including consulting companies as the “big four,” as well as companies that specialize in audits
  • third-party audits: are conducted by independent organizations which have no contractual relationship to the developing organization. Such organizations include independent researchers, journalists, law firms, regulators, civil society, non-governmental organizations (NGOs), users, and affected communities.
We adopt the audit design considerations that Raji et al. [70] has developed by analyzing various audit studies to account for the range of possible algorithmic audits. These considerations include (1) target identification and audit scope; (2) auditor independence; (3) auditor privileges; (4) auditor professionalization and conduct standards; and (5) when to audit and post-audit actions. Given these considerations, five audit types can be distinguished: algorithmic risk assessment, conformity assessments, research and civil society audits, regulatory inspection and certification, and sociotechnical audits. The characteristics of these audits and their assignment to first-party, second-party, and third-party audits can be found in Table 1.
First-and second-party assessments and audits Third-party assessments and audits
Algorithmic risk assessment, Algorithmic impact assessment
Ethic-based auditing
Conformity assessment
Research and civil society algorithm audits, bias audits, adversarial audits
Regulatory inspection and certification
Algorithmic impact evaluations, sociotechnical audits
Example(s)
Case study algorithmic impact assessment for data access in a healthcare context [33], data protection impact assessment [11]
End-to-end framework for internal algorithmic auditing by Raji et al. [69]
AIA conformity assessments, example procedure by [28]; Medical Device Regulation Application Procedure [16]
Gender Shades [13], Audit of the Rotterdams’ welfare fraud ADM [17, 18], Audit Austrian AMS algorithm [5]
UK Information Commissioner’s Office’s ‘Guidance on the AI auditing framework’[41]
Stanford’s Impact evaluation of a predictive risk modeling tool [31], sociotechnical audit that assess police use of facial recognition[66]
Auditors
Creators or commissioners of the algorithmic system, contractors such as consulting companies
Creators or commissioners of the algorithmic system, contractors such as consulting companies
Creators or commissioners of the algorithmic system, contractors such as consulting companies
Researchers, investigative journalists, data scientists, NGOs, affected communities, or other stakeholders
Regulator or other auditing and compliance professionals that have an institutional role
Researchers, policymakers, investigative journalists, data scientists, NGOs, affected communities, or other stakeholders
Target Identification & Audit Scope
Broad scope: focus on the risks and negative consequences of the deployment of an algorithmic system
Narrow scope: focus predominantly on technical prevention and uncovering of specific harms (e.g. bias audit)
Broad scope: focus on an algorithmic system’s compliance with regulation and mostly non-technical and process-oriented
Narrow scope: focus predominantly on technical prevention and uncovering of specific harms (e.g. bias audit)
Broad scope: focus on an algorithmic system’s compliance with regulation
Broad scope: assessing possible societal impacts of an algorithmic system on the users or population it affects after it is in use
Auditor Independence
Internal and thus no independence or contractual dependence
Internal and thus no independence or contractual dependence
Internal and thus no independence or contractual dependence
External by independent organizations with no obligation
External by independent institution, non-genuine independence when auditing administrative state-run ADMs
External by independent organizations with no obligation
Auditor Professionalization
Formal: done by deployers, internal teams or accredited second-parties
Formal: Done by deployers, internal teams or accredited second-parties
Formal: done by deployers, internal teams or accredited second-parties
Mostly informal: accreditation and cooperation with company possible , although basic access could be possible without accreditation (’right to scrape’)
Formal: Accreditation
Mostly informal: accreditation and cooperation with company possible (vetted researchers in DSA)
When to audit
Before deployment
Before deployment
Before deployment, post-market monitoring is possible
After deployment
After or before deployment, potentially an ongoing process
After deployment
Table 1: This table is inspired by and derived from [2, 53, 57, 70] and should give an overview of assessments and audits that we surveyed. The considerations of Raji et al. [70] were used to display the differences of these assessment and audit types. However, this table claims neither completeness nor complete adequacy, as this topic is a relatively emerging one with respect to algorithmic systems. At the same time, a systematic presentation of assessments and audits differences is not the focus of this paper, even though we conceptually argue and advocate a broad deployment of assessments and audits in the AI lifecycle with the AI audit ecosystem.
Algorithms cannot be divorced from the contexts in which they are applied [90] and represent a complex network containing economic, environmental, political, and social interests and outcomes affecting their functioning. The entanglement of technical components, and social elements leads to the typical acknowledgment of AI systems as sociotechnical systems [^4] Due to a broader understanding of their ecological impacts in recent years, [71] has added an “ecological” element to this notion, meaning that AI systems are socioecological-technical systems. Thus, it may be necessary to include ecological harms that AI causes in audits. [77] And, despite the common perception of audits as predominantly technical (see e.g. Koshiyama et al. [44]), internally conducted impact assessment as well as externally conducted impact evaluations and sociotechnical audits encompass a broader spectrum. They evaluate the impact of an algorithmic system on a population in a comprehensive framework that encompasses technical elements – most of the time limited technical elements – and include the social elements of the system [2]. We will come back to the assessment of sociotechnical systems in section 4.4.1.
Ethics-based internal auditing and adversarial audits by researchers and civil society have a more technical scope and tend to focus on specific harms, e.g. bias audits. Nevertheless, drawing a clear line of demarcation between narrow and broad scope of audit is difficult. While maintaining a however central focus on sociotechnical systems, it is essential to consider the data that substantiate claims concerning specific harms attributed to an AI system. An example of this difficult distinction is that Stahl et al. [82] includes bias mitigation assessments in the systematic review on impact evaluations, which we have introduced as ethics-based audits or bias audits.
Conformity assessments [56] and regulatory inspection [2] are systematic tests that evaluate if a system is compliant with a specification, standards, or a regulation. Regulatory inspections - conducted by an independent state-run body - are external to the deploying organization. Conformity assessments can be internal or external. In contrast to third-party audits, conformity assessments aim to achieve compliance, while third-party audits by researchers and civil society aim to prevent potential harm or uncover it by systematically evaluating the systems and checking the documents.
The vast number of actors and contexts in which AI systems function produces multiple complexities for their auditing, and this has affected the precision of related legal provisions. This applies especially when we try to answer questions such as, “What are we auditing for? Which harms count, and how are they defined?” [39]. This situation shapes the information we need to assess the system’s compatibility with a legal regime or ethical parameter, considering the contextual character of transparency and access to information [6].
We focus on third-party audits, especially those by researchers, academics, and journalists, as well as broader sociotechnical audits. Such audits are characterized by being conducted post-deployment through systematic evaluations, typically with little access to and independent of the organization deploying the algorithm [53, 19]. Thus, they have no formal contractual relationship with the audit target [67].
In recent years, there has been increasing public and academic concern about the influence of social media on public discourse, democratic processes, and institutions [7, 65]. Over the years, researchers have encountered significant challenges when attempting to access online platforms’ data for the purpose of studying their operations, potential risks, and impacts [62, 61]. Moreover, access to AI technologies beyond ML-driven social media recommendation systems, such as ADMs or general-purpose AI, presents a significant challenge [39]. Scientists and civil society are dependent on the voluntary granting of access, and sometimes only process access is granted, i.e. access to documentation (e.g Allhutter et al. [5]). Are we repeating the same mistake by granting large companies unwarranted authority without enabling civil society to thoroughly examine and understand their products for potential harm?
In response to these challenges, the DSA and the AIA have introduced distinct frameworks for audits, third-party audits and access to these technologies by research and civil society. In the following section, we will introduce and explore both of these frameworks, shedding light on how they address (third-party) audits of platforms and AI products.
3 The Regulatory Framework Based on the DSA and AIA on Third-Party Audits and Data Access
The goal of this chapter is to understand how the DSA and AIA regulate third-party audits and third-party data and model access. Therefore we will analyze the the DSA on audits and data access first. Subsequently, we will do the same for the current draft of the AIA. It should be noted that the following explanations refer to the draft of the European Parliament’s proposal.
3.1 Digital Services Act
The DSA aims to create a harmonized legal framework that is applicable to all online intermediary services provided in the EU. The regulation helps promote a safer digital space by regulating the distribution of (illegal) content. To increase accountability to the public, the regulation includes a number of specific obligations for providers of VLOPs and VLOSEs [23], including provisions such as specific transparency obligations on content moderation or the creation and publication of reports on risk assessments and mitigation procedures.
3.1.1 The Regulations on Independent Auditing in the DSA, Art. 37 DSA
In order to ensure compliance with the DSA, a “novel institutional ecosystem” [46] will be established that involves independent third parties in an additional oversight role. [^5] Cf. Recital No. 92 DSA. [27] Therefore, the officially promulgated version of the DSA asserts that VLOPs and VLOSEs are legally required to undergo an independent audit at least once a year (“independent audit”) [^6] Art. 37 (1) DSA. The purpose of auditing is to determine whether the providers are following the obligations standardized by the DSA. [^7] It means the obligations set out in Chapter III (Art. 11-48 DSA), in particular the obligations arising from codes of conduct referred to in Art. 45 and 46 DSA and the crisis protocols referred to in Art. 48 DSA. This is an important regulatory requirement, as it increases providers’ accountability and enables regulators and the public to understand and regulate how VLOPs and VLOSEs moderate content. The Commission is empowered to implement the requirements of Art. 37 DSA by legislative act. It has published a first draft of a regulation to specify the requirements for independent audits (hereinafter referred to as the Delegated Regulation on independent audits) [24]. The introduction of such guidelines is welcome as it will provide legal certainty for both audit firms and VLOPs and VLOSEs by providing guidance that can improve transparency and accountability. An important audit tool is the preparation of an audit report, which should comply with a comprehensive set of obligations. The audit firm must describe the specific elements audited and the methodology used. [^8] Art. 37 (4)(2d) DSA. It must also provide a description and summary of the main audit findings. [^9] Art. 37 (4)(2e) DSA. In addition, opinion must be given as to whether the audited provider has complied with the obligations and commitments referred to above. [^10] Art. 37 (4)(2g) DSA. The audit opinion shall state the result, which can be “positive”111111“Positive” means, where the auditing organisation concludes with a reasonable level of assurance that the audited provider has complied with an audited obligation or commitment, Art. 8 (1)(a) Delegated Regulation on Independent Audits [24]., “positive with comments”121212“Positive with comments” means, where the auditing organisation concludes with a reasonable level of assurance that the audited provider has complied with an audited obligation or commitment, Art. 8 (1)(b) Delegated Regulation on Independent Audits. or “negative”131313“Negative” means, where the auditing organisation concludes with a reasonable level of assurance that the audited provider has not complied with an audited obligation or commitment, Art. 8 (1)(c) Delegated Regulation on Independent Audits. In the case of a negative remark, the audit firm must make operational recommendations to be implemented within a recommended period of time to achieve compliance. [^14] Art. 37 (4)(2h) DSA If the audit company cannot conduct the auditing at all or can only do so partially, this should be clarified. [^15] Art. 37 (5) DSA. In such cases the Delegated Regulation on independent audits could specify how the systematic assessment of risks must be conducted. It should also specify the factors to be taken into account in the risk analysis. [^16] Cf. Art. 9 (4) Delegated Regulation on independent audits. According to the current draft the inherent risks171717“Inherent risks” means the risk of non-compliance intrinsically related to the nature, the activity and the use of the audited service, as well as the context in which it operates, and the risk of non-compliance related to the nature of the audited obligation or commitment, Art. 2 (10) Delegated Regulation on Independent Audits., the control risk181818“Control risks” means the risk that a misstatement is not prevented, detected and corrected in a timely manner by means of the audited provider’s internal controls, Art. 2 (11) Delegated Regulation on Independent Audits, and the detection risks191919“Detection risk” means the threshold beyond which deviations or misstatements by the audited provider, individually or aggregated, would reasonably affect the audit findings, conclusions and opinions, Art. 2 (12) Delegated Regulation on Independent Audits. should accounted for in the audit companies’ analysis. [^20] Art. 9 (3) Delegated Regulation on Independent Audits. Providers should provide the necessary cooperation and assistance to organizations conducting audits under this obligation so that they can undertake their work in an effective, efficient, and timely manner. [^21] Art. 37 (2)(1) DSA. They must, for example, grant access to all relevant data and premises. This includes the disclosure of data relating to algorithmic systems and answering written or oral questions. [^22] Cf. Recital No. 92 DSA. In addition to this, the Delegated Regulation on independent audits could regulate which information the providers have to provide to the audit organisations, Art. 5 (1) Delegated Regulation on Independent Audits. The provider and the audit organisation must necessarily agree on conditions - i.e. meaning duties and obligations - that are listed in the current draft to carry out the audit, Art. 7 (1)(a) Delegated Regulation on Independent Audits. The platform provider must then consider improvement suggestions and take action to enforce the recommendations within one month. [^23] Art. 37 (6)(2) DSA. If the operational recommendations are not implemented, the VLOP or VLOSE provider must justify and specify alternative measures that will be taken to fix the identified cases of non-compliance. [^24] Art. 37 (6)(3) DSA.
3.1.2 The Requirements for Audit Firms under the DSA
The auditors should have no conflicts of interest with the provider. [^25] Art. 37 (3a) DSA. In this respect, the European legislator clarifies that audit firms are organizationally independent from the interests of the provider and constitute an external audit structure. The additions were explicitly included to address concerns that audit firms are dependent on platform providers for auditing. Cf. critical comments on the preliminary draft [14]. Furthermore, they must have proven expertise in risk management as well as technical competence and capability.  [^26] Art. 37 (3b) DSA. This requires that they must have the necessary knowledge in the field of risk management as well as the technical expertise to test algorithms, cf. Recital No. 92 DSA. Providers are allowed to choose the audit firms themselves. However, it is important to note that the DSA introduced a rotation model according to which the audit firms may only work for the providers if they have not performed an audit pursuant to Art. 37 DSA for a period of more than ten consecutive years. [^27] Cf. Art. 37 (3a)(ii) DSA.
3.1.3 Data Access for Vetted Researchers, Art. 40 DSA
Art. 40 DSA regulates access to the data of VLOPs and VLOSEs for “vetted researchers”.The norm contains not only the obligation of providers to grant access, but also the right of researchers to access.  [^28] E. g. [38] and [91] Researchers must undertake to make their research results publicly available free of charge  [^29] Art. 40 (8)(1g) DSA.. The results can thus be made available to the authorities, the Commission and the public. Classification as an approved researcher is subject to strict conditions. In particular, researchers must be affiliated to a research institution  [^30] Art. 40 (8)(1a) DSA.. It means a university, including its libraries, a research institute or any other entity, the primary goal of which is to conduct scientific research or to carry out educational activities involving also the conduct of scientific research.  [^31] Art. 2 (1) Directive (EU) 2019/790. Like the audit firms, the researchers and the institutions to which they belong must be independent of commercial interests  [^32] Art. 40 (8)(1b) DSA.. Another condition for access is that the data collection is necessary and proportionate for the creation of the data work.  [^33] Art. 40 (8)(1e) DSA. Cf. Recital 97. The classification as vetted researcher is granted to the researcher only in relation to the intended research project. The research must relate to systemic risks and contribute to their identification and understanding. [^34] Art. 40 (4) and (8f) DSA.
3.2 The Artificial Intelligence Act
The AIA [22, 26] aims to regulate and harmonise the marketing, commissioning and use of AI systems within the European Union. To this end, safe and trustworthy AI is to be created by observing the fundamental values protected in the European Union, that is material and ethical values, during development and use. For this reason, companies [^35] The regulation is addressed to all companies that place AI systems on the market or put them into operation in the European Union, Art. 2 (1) AIA [22, 26]. In practice, this will include software developers, AI importers and distributors, but also, for example, government authorities. should be subject to a far-reaching catalogue of obligations and ensure quality assurance throughout the entire AI value chain. In contrast to the Commission, the European Parliament’s draft also takes account of foundation models, which would be subject to specific regulatory requirements [96].
3.2.1 The Regulations on the Conformity Assessment Procedure
Basically, in the context of conformity assessment procedures, the AIA focuses primarily on internal control by the first parties. [^36] Annex VI AIA. The AIA opens up the possibility of having such procedures performed by external companies only for so-called high-risk systems [^37] Cf. Art. 6 AIA and Annex III for a description of which AI systems could be high-risk system.. [^38] Art. 43 (1)(1b) AIA. However, whether a provider of an AI system uses a notified body depends on the nature of the high-risk system. According to the current draft, it remains optional if, for example, biometric and biometrics-based systems are used, which are partly subject to harmonised standards that cover the relevant security requirements for the AI system in the AIA and common specifications exist in accordance with Art. 41 AIA. [^39] Art. 43 (1)(subpar. 1) AIA: “[…] the provider shall opt for one of the following procedures:”. On the other hand, the conformity assessment procedure by external bodies is mandatory, for example, when harmonised standards are only partially applicable to the AI system or if the provider considers an external audit to be necessary. [^40] Cf. Art. 43 (1)(subpar. 2) AIA. An audit should also be carried out if the AI system is already subject to a conformity assessment procedure, but there has been a significant change to the system, or the change is as a result of the system’s learning process. [^41] Cf. Art. 43 (4) AIA. In the other cases, [^42] This refers to high-risk systems in the sense of Annex III Nr. 2-8 AIA. the AI systems are only subject to internal control. They are not subject to conformity assessment procedure within the meaning of the AIA [^43] Cf. Art. 43 (2)(1) AIA. or a conformity assessment procedure according to other regulations under European law. [^44] Art. 43 (3)(subpar. 1)(1) AIA. In the latter case, however, the provider of the AI system must take into account some of the procedural requirements of the AIA. [^45] Art. 43 (3)(subpar. 1)(2) AIA. The requirements for the conformity assessment procedure are set out in Annex VII. The procedure involving a notified body should be conducted in a way that preserves the confidentiality of the information provided. [^46] Cf. Art. 33 (6) and 70 (1) AIA. It is composed of the audit and monitoring of the respective approved quality management system, [^47] Annex VII No. 3, No. 5 AIA as well as the control of the technical documentation obligations [^48] Annex VII No. 4 AIA. imposed on the providers. Furthermore, the audit firm may be granted access to the training models and trained models of the AI system, including its relevant parameters upon request, and they should be able to test the AI system independently. [^49] Annex VII No. 4.4 and No. 4.5 AIA Provided that the notified body can confirm the conformity of the high-risk system, a certificate — similar to the audit opinion under the DSA — should be issued. [^50] Art. 44 (1) and (2) AIA. That certificate would have a duration of up to five years or could be revoked or limited if the AI system later fails to meet the requirements of the AIA. [^51] Art. 44 (3) AIA. The providers of the AI system should support the audit organizations in their investigations by, for example, providing the technical documentation that makes available all necessary information to assess whether the AI system complies with these requirements. [^52] Art. 11 (1) AIA. Non-compliance with the support services is subject to a fine of up to EUR 5 million or 1 per cent of total annual worldwide turnover, Art. 71 (5) AIA. They have to grant the audit company unrestricted access to the training and test data sets that the provider used. This also applies to application programming interfaces and other means and tools suitable for remote access. [^53] Annex VII No. 4.3 AIA. In order to adapt the regulatory framework, the power to adopt acts amending the provisions on conformity assessment procedures should be delegated to the Commission. [^54] Recital No 85 (1) AIA.
3.2.2 The Requirements for Audit Firms under the AIA
So-called “notified bodies” should perform the conformity assessment procedures. [^55] Art. 33 (1) AIA. A “notified body” is defined in the AIA as a conformity assessment body notified in accordance with the AIA and other relevant Union harmonisation legislation. [^56] Art. 3 (1)(22) AIA. The competent authority [^57] This is the national authority of the respective Member State to which the monitoring of compliance with the AI Act has been assigned. According to Art. 3 (1)(19) AIA, a designated authority is the national authority responsible for establishing and carrying out the necessary procedures for the assessment, designation, and notification of conformity assessment bodies and for their monitoring. Among the tasks of the notifying authority, cf. Art. 59 AIA. then notifies them, provided that they fulfil the specific requirements in the AIA. [^58] Art. 31 (1), 32 (1) AIA. The procedure for designating qualified bodies specifies requirements for (professional) qualifications as well as their independence and objectivity. As in the DSA, the AIA also clarifies that the notified bodies should be independent of the provider of the high-risk system. [^59] Art. 33 (4) and (5) AIA. Cf. also the European Parliament’s clarification that the assessment may only be carried out by staff who have neither worked for the high-risk scheme provider nor been employed in a related legal entity for twelve months prior to the assessment, Art. 33 (4a) AIA. In this context, the independence of the notified body extends to the provider of an assessed AI system as well as to all other actors with an economic interest in thas AI system. [^60] Cf. Art. 33 (4) AIA. In order to ensure professional competence, they must meet the specific organisational, quality management, resourcing, and procedural requirements necessary for the conformity assessment of AI systems. [^61] Cf. Art. 33 (2), (3), (10) AIA. To this end, the notified bodies should have procedures guiding their activities which take due account of the size of an undertaking, the sector in which it operates, its structure and the degree of complexity of the AI system concerned. [^62] Cf. Art. 33 (9) AIA. In principle, the provider of the AI system can choose between the notified bodies, unless the AI system is put into operation by a law enforcement, immigration or asylum authority or by EU institutions, bodies or agencies. [^63] Cf. Art. 43 (1)(subpar. 3) AIA. Then the market surveillance authority mentioned in Art. 63 (5) or (6) AIA takes over the function of the notified body. In contrast to the DSA, the AIA ensures close cooperation with the notifying authority. The state participation control is enforced by the certification of the bodies to be inspected. Furthermore, the certification bodies are accountable to the notifying authorities. In order to qualify as a notified body in the sense of the AIA, the potential audit company has to fulfil a number of requirements: Extensive information must be added to the application for notification, such as the conformity assessment activities, the conformity assessment modules and the AI technologies for which the conformity assessment body claims competence, and – if available – an accreditation certificate. [^64] The accreditation certificate is issued by a national accreditation body and certifies that the conformity assessment body complies with the requirements of the AIA, Art. 31 (2) AIA. In the absence of a corresponding accreditation certificate, the necessary evidence for the verification, recognition and periodic monitoring of compliance with the requirements of Art. 33 AIA shall be sent to the designated authority in accordance with AIA[56].
3.2.3 No Data Access for Vetted Researchers
The AIA does not contain a regulation comparable to Art. 40 DSA. This is surprising, as the regulation aims to ensure that relevant entities, such as digital innovation hubs, testing experimentation facilities and researchers, should be able to access and use high quality datasets within their respective fields of activities for the development and assessment of high-risk AI systems. [^65] Recital No. 45 (1) AIA. The European data rooms created by the AIA and the facilitated exchange of data should also provide trustful, accountable and non-discriminatory access to high quality data for the training, validation and testing of AI systems. [^66] Recital No. 45 (2) AIA.
4 Creating a Diverse AI Audit Ecosystem for Accountability by Involving Research and Civil Society
In this section, we describe an AI audit ecosystem that ensures two key accountability components. For that, we draw on a definition of accountability in the AI context. We look at three aspects which demonstrate why internal audits are not sufficient to create an audit ecosystem. In three additional aspects, we point out why third-party audits are valuable for an AI audit ecosystem . This will be the blueprint under which we will explore whether the analyzed EU legislation can create an AI audit ecosystem.
4.1 Third-Party Audits by Researchers and Civil Society
[13, 74, 59, 45, 83] are some examples of algorithm audits in the United States that have uncovered problematic algorithmic behavior or algorithm impacts that could potentially cause harm. Regarding algorithmic audits in the EU, audits of the Austrian AMS algorithm [5], and Rotterdam’s welfare fraud algorithm by the Wire [17], come to mind. These audits were third-party audits, that is, audits by independent, external parties combining independent researchers, civil society, regulators, journalists, and other stakeholders. They have uncovered sociotechnical harms of algorithmic systems such as representational harms due to discriminatory choice.
With respect to social media platforms, third-party audits uncovered social system harms due to the propagation of disinformation on social media, hate speech increase on Twitter [37] and uncovered bias with respect to right-winged content recommendation on YouTube [74]. It has to be noted that third-party audits by researchers and civil society are not conventionally considered audits per se due to their narrow target specification [70], but they have repeatedly shown their significance. Our argument, therefore, favors their integration in an AI audit ecosystem.
Third-party audits aim for a material change in the situation (e.g., a product update, policy change, or recall) to minimize the harm experienced by those they represent [67]. In contrast, first- and second-party audits that seek to validate procedural expectations, aim to minimize liability and solely test for compliance to AI principles and legal constraints [67]. Despite the important oversight of third-party audits in platform research and algorithmic systems, it has been difficult for them to audit these models formally and legally. Until now, evaluations are usually done through the application developers’ voluntary public disclosures of possibly incomplete transparency reports or by voluntarily providing certain data access through an API [3].
4.1.1 Third-Party Platform Data Access
Olteanu et al. [62] reports four obstacles with systematic platform research through third parties: (1) Many social platforms discourage data collection by third parties as some platforms such as Facebook block their API for researchers and other platforms that block scrapers. Additionally, there is a shift toward APIs of major platforms and products being closed, a scenario described as the “post-API scenario” [86]. This leads to more open-source products or products being audited by third parties by deployers, who are probably more concerned about the compliance of their products. This leads ultimately to an imbalance in oversight [70]. (2) Even API access comes with limitations as they limit the quantity of data, increase the monthly costs and provide only query languages with limited expressiveness. (3) Additionally, these platforms may not give access to all data as they safeguard privacy-related content although the platforms could anonymize them. (4) Finally, the platform may revoke access every time.
For example, access for researchers has changed recently for the Twitter API [63] and the Reddit API, as both have raised the costs for third-party API access [73]. Nevertheless, attempts by third parties such as NGOs, research and investigative journalists to expose these accountability concerns by data donations or scraping are thwarted by companies. For example, the NGO AlgorithmWatch was prevented to conduct an audit of the Instagram algorithm via data donations. Facebook sued them for disregarding the terms of service [40]. Data donations are only necessary because the APIs for limited ’black-box’ access have been shut down or made considerably more expensive by many platforms and services [73, 63, 86].
The DSA will provide data access for vetted researchers. However, as demonstrated vetted researchers may not include NGOs and investigative journalists, although many of the case studies mentioned were conducted by newspapers and were accompanied by NGOs that were able to contribute with their domain knowledge to potential harms. We will discuss reasons for this being critical later.
4.1.2 Third-Party ADM and ML-Model Access
It appears to be just as challenging with ADMs and ML-driven products. Many of the products do not have access for researchers and regulatory audits. Therefore, mostly commercial APIs have to be used as in the Gender Shades case. If there is no API, the data can only be scraped, which is mostly illegal or prevented by other means. US Researchers are currently campaigning against the use of anti-hacking laws in the US that prevent audit studies by scraping. At the same time, according to Raji et al. [70], companies continue to seek to prevent audits by (1) paywalls, (2) prohibition laws through terms of services, and (3) structuring the product to obscure any clear set of test points. Raji et al. [70] describes it as an extraordinary effort for researchers, NGOs, investigative journalists to gain access and knowledge of ML and ADM systems.
It was shown that the AIA does not encompass data or model access provisions for third parties, even in the context of high-risk systems. Nonetheless, in the subsequent discussion, we will explore the potential access levels and, consequently, the knowledge accessible to third parties concerning AI technologies. Based on these findings, we will determine to what extent it is possible to make statements about a system even with limited access.
4.1.3 Technical Methods and Access Levels
Third-party audit methods [76, 8, 84, 54] – such as code audits, user audits, automated audits (including scraping audits, sock puppet audits and API audits), document audits, crowdsourced audits, and experimental audits [84] – contribute various approaches to auditing platforms and AI systems. Whether social media platform, ADM or ML model audits, the methods for third-party audits are similar.
Document audits are often broad sets of information that can analysed in different ways [54] and ensure the possiblity for third-party impact evaluations or conformity assessments as intended by the AIA. However, only process-based compliance is feasible through document audits, there are no standards for most of these reports and they are impossible to verify if no other data access is provided [54].
Code audits entail direct access to a system’s code and involve the dynamic execution of code with various inputs to observe corresponding output behaviors [84]. While they are suitable for some audits, code audits might not be ideal for AI product audits because the need for training data. Furthermore, potential complexities in reverse engineering system behavior - especially for dynamically trained algorithms such as recommender systems - hamper code audits’ feasibility.
Scraping audits, on the other hand, lack randomization and interventions. This limits their utility primarily to descriptive tasks [8]. They additionally lack personalization for specific user accounts [54]. Furthermore, scraping and sock puppet audits often violate terms of services of platforms [54]. API access limitations were already discussed.
Experimental audits involve the change of the algorithm and can thereby establish causal evidence since interventions are feasible [29]. However, full access to the system is necessary for experimental audits, and experiments with social platforms such as Facebook on emotional contagion exhibit ethical difficulties [78].
Crowdsourced audits, including data donations, have two major advantages: They include users in the audit, and they transfer parts of the training data so that more statements are possible. However, in the past, large platforms or companies have prevented such crowdsourcing audits as in the case of AlgorithmWatch. Recognizing the inherent advantages and limitations across various audit types, a comprehensive analysis of these factors would extend beyond the scope of this article [^67] For a comprehensive analysis of connections between audit types and recommender system elements see [54].. Consequently, our focus is confined to examining access to AI systems.
Technical access and the knowledge of a system can be categorized on a continuum from “white-box” to “black-box” and is related to these audit types. Koshiyama et al. [44] have defined seven levels of auditing access that range from “process-access,” where only indirect observation of a system is possible, and the system’s behavior in a sociotechnical system must be reverse-engineered, to “white-box,” where details of the encompassing model are disclosed. The latter is usually not applicable to third-party audits, as it would conceal all information about the model, which could have security, privacy, and trade secret issues [44]. Thus, Koshiyama et al. [44] have described an “information concealed versus feedback detail trade-off.”
What is accessed
Tools and possible evaluations
Feasible audit types
Case study third-party audit(s)
Process access
Frameworks, checklists, model Cards, datasheets, method paper, technical reports
Checklists for explainability, robustness, fairness and privacy
Document audits (algorithmic risk assessment and algorithmic impact evaluations)
Audit Austrian AMS algorithm[5]
Model access (’black-box’)
Access to the predictor model $f(\cdot)$ using artificial data $x^{\ast}$
Feature relevance plots, partial dependency plots, adversarial attacks, adversarial fairness, statistical disclosure
Scraping, API, sock puppet audit and code Audit
Bloomberg stable diffusion audit[59], discrimination in online ad delivery, [83], auditing radicalization pathways on YouTube [74]
Input access
Access to the predictor model $f(\cdot)$ and training data $x$ to predict $f(x)$
Surrogate explanations, synthetic data, bias in outcome
Crowdsourced Audit686868just a fraction of training data $x_{p}\in x$, API audit with training data
Cooperative audit of pymetrics’ candidate screening software[93]
Outcome access (’grey-box’)
Access to the predictor model $f(\cdot)$, training data $x$ to predict $f(x)=\hat{y}$ and the actual outcome $y$
Accuracy of explanations, concept drift analysis, bias in opportunity, inversion attacks
API audit with training data and history of actual outcomes
ProRepublicas investigation of North- pointe’s recidivism risk scoring system [45]
Parameter control and learning goal access
Learning procedure and target $L$, parameters $\phi$, $L(f_{\phi}(x))$ can be retrained
Stability of explanations, stability analysis, stability of bias metrics, functionality stealing, model complexity, stress-testing, trade-off of bias and loss metric, model extraction
Retraining of the actual model with different parameters
Audit of the Rotterdams’ welfare fraud ADM[17, 18]
white-box
Access to the whole architecture $f$ and all documents
Documents and specific explanations, model selection and validation, model selection and development, model security evaluation
Experimental audits (interventions)
Accessible by first-parties and possibly second-parties
Table 2: The Table presents leveled access modes from ’process access’ to ’white-box access’. We connected access modes with ’what is accessed’, tools and feasible audit types. The table must be read so that in each access mode possible tools, audit types and accesses are added to those that were in the level before. This may not be the case only for process access, since documents are not available in each access level. The table is inspired and derived from Koshiyama et al. [44], Koshiyama et al. [43] and Brown et al. [12]. As in the case of table 1, this table should provide only an overview. We derived the table via literature review and not via systematic methods. A systematic evaluation is probably not conductable as the field is still emerging. However, we want to show how data access and audit methods are connected and demonstrate the trade-off between access and revealed information.
Table 2 shows different access levels that Koshiyama et al. [44] have proposed for ML algorithm audits. We matched them with third-party evaluation tools, feasible audit types, and case study third-party audits. This representation is a simplification because platforms and general-purpose AI consist of several models that interact with each other and ADMs can consist of several models and a complex decision function. Nevertheless, the presentation is useful to show three aspects. First, with more access, more knowledge about a system is possible and more audit types are available. Second, there is a trade-off between trade secrets and model access. Third, case study third-party audits in the past that had only ’black-box’ access - mostly through an API, scraping, or sock puppets - demonstrate that low-level access can contribute to accountability.
Some audits, such as the Austrian AMS algorithm relied solely on process access. The prediction model has not been made public [5]. Generally, releasing personal data relating to ADMs would obviously be dangerous. In this case, anonymization methods would have to be used to release the data for scientific purposes like it is planned by the DSA. These document audits have little expressiveness as they have to reverse engineer the whole system’s functioning without the knowledge or verification that the documents are actually accurate. Others with at least ’black-box’ access could provide more information.
For example, Bloomberg’s stable diffusion audit [59], or in YouTube’s radicalization audit [74], the prediction model was accessed via the API. Thus, problematic behaviour like biases could already be shown in both cases. The original gender shades study as well operated within the constraints of accessing their audit targets via a commercial API, mirroring the actions of a user within these platforms and thus using a socket puppet [68]. Others had even more access. In the case of the cooperative audit of pymetrics’ candidate screening software the auditors had ’grey-box’ access, as the outcome is not feasible [^69] As it’s unknowable if someone would have performed well that didn’t get the job. for recruitment algorithms [^70] This audit was criticized as [95] claimed that the authors got funded by the company for reducing the scope of the audit. If this claim is justified, this would mean that the audit would not been conducted by independent external auditors, but dependent and thus internal auditors. The example again demonstrates the power of tech companies.[94]. In the case of ProRepublicas investigation of the recidivism risk scoring system they had access to the model, the training data and the actual outcome [45]. In the case of Rotterdam’s welfare fraud audit, the auditors had parameter control and learning goal access [17]. Model access alone has shown its effectiveness through certain case study audits. Audits, including Bloomberg’s stable diffusion audit [59], examinations of discrimination in online ad delivery [83], and the audit of radicalization pathways on YouTube [74], have achieved auditing outcomes with limited access, albeit frequently encountering challenges due to legal constraints. Especially in the case of gender shades, where after the audit there was an improvement in several facial recognition tools in terms of functionality for black women [68]. This proofs limited ’black-box’ access for researchers can contribute to mitigating harm.
While other levels of access might provide more comprehensive insights into a system, the prevalence of reverse-engineering methods underscores the significant influence of third-party audits. Admittedly, a system-tailored approach is necessary, given the nuances involved, and model-agnostic audits, such as partial dependency plots and feature relevance plots, present challenges [35]. Nonetheless, it is evident that third-party audits, even with ’black-box’ access, have managed to provide insights when combined with evaluations of the sociotechnical framework. Particularly because they contribute to a system for which the depth of knowledge acquired corresponds to the access level, audits with even minimal access can generate momentum for obtaining greater access. In some cases, this may involve collaborating with other auditors and stakeholders. This concept is further explored in the next chapter, “AI Audit Ecosystem.” It is worth noting that substantial examination by sock puppets or API audits is often feasible only for products that possess open APIs.
4.2 An AI Audit Ecosystem
Algorithmic accountability regards a networked account for an algorithmic system, where several actors must explain and justify to a forum or several fora their use or design of an algorithmic system, as well as the decisions and the subsequent effects concerning the system . The fora can be internal and external, as well as formal and informal [92]. Novelli et al. [60] have defined the goals of AI accountability as (1) compliance, (2) reporting, (3) oversight, and (4) enforcement. Although audits interfere with reporting and enforcement, they primarily achieve compliance and oversight. Compliance concerns the binding of the system to align with ethical, legal, or technical norms. For example, performing an ethical assessment may constitute compliance. Typically, compliance refers to compliance with the law. Oversight seeks to find relevant facts about the system, such as performance, bias metrics, or other measures, through an overseeing body that can be internal or external.
In order to ensure oversight and compliance, Raji et al. [70] and Albert [3] have proposed establishing an AI audit ecosystem that includes internal and external auditors. Sambasivan et al. [75] have called for the establishment of an ecosystem for accountability in India that includes various stakeholders and empowers oppressed communities. Mökander et al. [56], meanwhile, has argued that the AIA de facto proposes to establish an AI audit ecosystem. However, none of these authors have described what constitutes an audit ecosystem or why we need one. While following this line of reasoning about the importance of including diverse auditors, we define an audit ecosystem and argue that the AIA does not establish an AI audit ecosystem.
We define an AI audit ecosystem as an open dynamic partial self-organizing community of hierarchically independent, yet interdependent heterogeneous auditors and audit types [^71] Inspired by ecological, digital and innovation ecosystems: ”a community of hierarchically independent, yet interdependent heterogeneous participants who collectively generate an ecosystem output”[85].. The characteristics are described by the following:
  • open, meaning that it interferes with other accountability processes, such as reporting, enforcement, and harm incident reporting
  • dynamic over time (and thereby adaptive to new forms of AI systems, which may have to be audited differently) and over stakeholders (it must involve potentially affected communities and users)
  • partial self-organization of auditors in an enabling environment [^72] See in anti-corruption social accountability ecosystems [34] and a promotion of sharing audit information through interaction between auditors (formal and informal)
The audit ecosystem is needed to ensure compliance and oversight throughout the AI lifecycle. We argue that for such an ecosystem, third-party audits by researchers and civil society are a vital component to fulfill the characteristics of the ecosystem and thus also provide both oversight and compliance through first-party audits. Novelli et al. [60] have stated that ”[o]verseers may act at different levels that may overlap, e.g., an internal audit is compatible with judicial review accountability.” However, we argue that overseers should act at different levels, as (1) internal audits are prone to audit-washing as well as false assurances, and they may only ensure compliance (see section 4.3.1) and are (2) not adaptive to new technologies (see section 4.3.3). Additionally we argue that (3) NGOs and researchers possess valuable experience in incorporating a holistic perspective that encompasses socioecological-technical systems. They are able to draw from a background in conducting impact assessments and human rights considerations (see section 4.4.1) and (4) tend to include affected communities more than companies as they have other audience and incentives (see section 4.4.2).
4.3 Why Are Internal Audits Insufficient?
Internal audits and assessments have emerged as a prevalent focus in the academic discourse. In alignment with this trend several regulations – the Algorithmic Accountability Act of 2022 in the US, the AIA, and GDPR enforcement – call for first-party audits, underscoring their outsized focus in academia and governance [70]. Internal audits have the advantage that the developers know their system best and have “white-box” access [44]. This provides them the most technical knowledge about the system and they can intervene at any stage of the design process as well as throughout the AI lifecycle. It also puts the burden of responsibility for harm prevention on the shoulders of the developers and deployers  [^73] If both parties are the same company. and thus on those who establish a business model with it. Nonetheless, as demonstrated by precedent audit domains, the adequacy of internal audits in ensuring comprehensive oversight has been called into question, as [70] has highlighted.
4.3.1 “Audit Washing” and False Assurances
Furthermore, the overreliance on internal audits could lead to unverifiable assertions that the AI systems has passed ethical and legal standards, leading to even more harm and less oversight[32]. There is an ongoing discussion within research communities on how to translate ethics into practice. The discussion evolves around which fairness metrics should be used while as fairness remains a contested term [55]. General standards and norms for audits [19] are therefore absent. This leads to an environment conducive to the phenomenon of “audit-washing”. A firm’s business goals may not always align with harm-reduction through course-correction or ethical design of their products. Developers and deployers may ignore audit recommendations that threaten their business interests [58].
Empirical research substantiate these concerns. [52] have discovered that instructing developers to incorporate ethical codes yielded no notable change in their established practices. Moreover, studies involving AI development companies [88] and startup-like environments[87] have shown a disconnect between acknowledging the significance of ethics and integrating them into AI practices, indicating a substantial gap between ethical research and practical implementation and auditing.
Such absence of standards and norms in corporate environments could lead to using tools and metrics that play down the risks of ones product [39]. Statistically, through p-hacking or data dredging, and more generally through fairness gerrymandering, it may always be possible to suggest that your system is fair [8]. Issues concerning independence of second-party auditors were discussed frequently in financial audits. The opinion paper on this issue of Bazerman et al. [9] highlights the the emerging bias towards the client in contractual relationship audits. In addition to intentional false assurances of the absence of harm, audits may be performed incorrectly or issues may go unaddressed. That is what happened in the Gender Shades audit. It revealed that the U.S. National Institute of Technology (NIST) failed to measure demographic differences in their regular Facial Recognition Vendor’s Test [70].
Several possible adjustments address for these issues. Costanza-Chock et al. [19] have proposed mandating (1) the publishing of (first-party) audit results, (2) harm incident reporting and other measures to detect and report harms, (3) the involvement of affected parties, and (4) the consideration of accreditation and standards for evaluation. Although all points are valuable for an audit ecosystem, we argue that the involvement of affected parties and standards for evaluation are only valuable if third-party audits support the ecosystem. (1) and (2) are particularly important for an open and balanced ecosystem, as information-sharing is promoted, and multiple self-organized auditors can immediately note and review potential harm incidents. While (4) is especially important for functioning enforcement, (3) is a basic condition for a dynamic audit ecosystem; otherwise, possible harm remains unnoticed.
4.3.2 Standards, Gaming the System and Goodhart’s Law
The development of standards is vital for compliance and verification. However, hard standards and technical targets (e.g. bias metrics) can also lead to decoupling from the intended purpose. In financial audits, decoupling was described as an audit process that becomes ”ceremonial” and rote. This can happen when the entity undergoing auditing has internal control structures that ritualize and channel outside audits [64].
Gaming the system is connected to Goodhart’s law, which is usually paraphrased as ”When a measure becomes a target, it ceases to be a good measure.” Thus, behaviors are created to achieve the metrics while pursuing other business incentives simultaneously. It could encourage adherence to metrics while maintaining business. Methods to game the system emerge to avoid implementing the desired behavior measured by the metrics [64]. For example, in the financial crisis, banks went bankrupt, which had been recently testified solvent [80].
The Wirecard scandal in Germany shows that hard standards do not necessarily help to uncover harm. Even though the company was audited several times by different second-party auditors it was still not revealed that the business model of Wirecard was a financial fraud [49]. In both cases, the companies learned to game the system, and game the standards. The Wirecard scandal was ultimately uncovered by investigative journalists [51]. This example demonstrates once again that there should be external oversight by such institutions, which do not have to adhere to hard standards and entrenched audit processes. This is all the more evident when it comes to new technologies that are developing rapidly.
4.3.3 Non-adaptivity to New AI Technologies Through Hard Standards
In an audit system of hard standards, ’gaming the system’ can occur from companies trying to exploit those standards and find loopholes. At the same time, new technologies in particular can create standardization gaps. To create an adaptive and dynamic audit ecosystem informal audits without hard standards and processes could help to find harm.
Social accountability in the public sector has demonstrated that strict process-based evaluations often slow down government action and suppress creativity. Therefore performance-based evaluations in audits are preferable to process-based ones. The role of civil society can be crucial in that regard, as it functions as a vigilant watchdog over both procedural integrity and performance outcomes. The involvement of civil society can introduce innovative performance metrics to supplement existing measures [1].
This becomes particularly evident if there is no external oversight from third parties capable of reproducing these evaluations. Novel AI products and technologies could necessitate the development of new evaluation methods. For example, in the aforementioned Gender Shades case or in recent examples of LLMs and stable diffusion. Although the relevant legislation and standards do not yet exist, it is difficult to imagine that if evaluation standards had existed, they would have worked for these type of AI systems. Developing standards for evaluating technology will take a long time, which is why it is necessary that third parties with limited access to the model engage in evaluating and understanding emerging systems. This happened with the stable diffusion audit by Bloomberg[59]  [^74] Reporters used stable diffusion to generate 5,100 images of people using a simple recurring prompt iterating through different categories. Through data-labeling and with ChatGPT, which researchers extensively evaluated and which showed biases [72].
4.4 Why Should We Include Research and Civil Society in the AI Audit Ecosystem?
4.4.1 Third-party Domain Knowledge
We emphasized the significance of auditing both the visible external connections and the internal self-regulatory mechanisms of algorithmic systems. NGOs, researchers, journalists, and members of civil society offer domain expertise that is absent in the organizations deploying algorithms [1]. Identifying such contextualities involves a broad range of perspectives about the impacts that such systems might entail. AI experts might not be capable of addressing this issue, and sometimes they are less capable.
According to Li et al. [48], they are not always aware of the way digital technologies affect marginalized communities, and their tools for investigating bias, for example, might not necessarily detect behavior that could not be reasonably predicted before its deployment. Instead, the authors write, “most expert-led auditing methods were developed to detect statistical disparities not, for example, if an algorithm is censuring or harmfully depicting a marginalized community in images or the provision of online services.” This is why we argue that these third parties potentially have specific domain knowledge that include sociotechnical system thinking, as they are have knowledge of human rights impacts in specific domains and are closer to marginalized or affected communities, which brings us to the next aspect.
4.4.2 Inclusion of Affected Communities
Pro-accountability endeavors that encompass diverse interests and ideologies incorporate more legitimacy than those driven by a limited group of professionals [1]. This is why scholars (e.g. [19], [48]) have recommended the inclusion of stakeholders who are or can be affected by AI technologies, or their representatives, in audits.
We call this a participatory audit [^75] At this point, it is important to note that the terminology surrounding what we call participatory audits has not yet been consolidated. When affirming the relevance of participatory audits in the context of governments’ accountability,[10] refers to “social audits” in a way that is very similar to what we describe as participatory. According to him, they differ from other forms of audits because they are performed by nonexperts and rely on engagement from citizens and/or civil society organizations. It is a mechanism of oversight, of exerting control on government officials to ensure transparency, responsiveness, and effectiveness by putting citizens in the position of active participants, not mere sources of information [10]., as it consists of a framework that aims to promote the participation of affected stakeholders in different aspects and moments of the design, development, and deployment of AI systems. Audits that consider individual or collective accounts of experiences related to algorithmic bias, harm, or injustice through incident reporting are currently predominantly third-party algorithmic audits [71].
Some authors have cited similar approaches for AI auditing as user-engaged approaches; these would consist of the direct participation of users of AI-based applications in surfacing harmful algorithmic behaviors in activities called “bias bounty” [30], for instance. However, since those affected by AI systems’ supply chains will not always be users of these technologies, we do not consider this to include affected communities [^76] They can also be workers who train AI systems to identify images correctly or in the case of ADMs, where decisions are made about affected people. It is possible that some of these individuals will never use these tools.. In this sense, one can frame participatory audits as a set of different methodologies to include users or affected individuals and groups as a whole in AI auditing processes.
Through methods such as workshops, bias bounty programs, and direct interviews, individuals potentially harmed by these systems can identify problems and help find solutions. NGOs, civil society organizations, and investigative journalists have the opportunity to involve affected communities and users. Of course, companies may have the capability to do the same; however, this appears to be highly improbable with regard to the company’s incentives. Successful community engagement depends on the willingness of the audited organization to genuinely listen, respond, and act on the concerns and recommendations identified during the audit process [1]. Affected communities may trust third-party auditors more than internal audits conducted by the organization itself. This trust can encourage community members to openly share their perspectives, concerns, and experiences during the audit process.
4.4.3 Information and Power Asymmetries
In the past, understanding and evaluating platforms has encountered considerable challenges due to insufficient transparency in their operations despite their vast power and influence. For years, users and regulators were frequently unaware of these digital platforms’ internal functioning, inhibiting the effective resolution of misinformation, data protection, and content moderation. Although the DSA points in the right direction by challenging information and power asymmetries, ’audit capture’ is possible. The risk is that large tech companies leverage their market power against their new mandatory auditors [46]. Because we have let the market power of large tech companies evolve for years without proper regulation.
A comparable scenario is occurring with AI products with the focus on conformity assessments by the AIA. In general power asymmetries are present in the standardization process, too. As the process of standardizing audits is ongoing, and the extent of public involvement remains uncertain, standards are usually developed by standard-setting organizations in which industry representatives exert significant influence. In contrast, the voices of civil society and consumer advocates are often overheard [15]. Therefore, researchers [20] demand that policymakers take the necessary steps to improve the overall standardization process, including the structural and organizational framework of standardization organizations, to facilitate an inclusive and democratic system that provides for broad stakeholder engagement and dialogue and input on the development of technical standards. Third-party data access is one measure to involve civil society.
We have seen in section 4.1.3 that third-party audits can uncover harms despite limited access. To understand what harms are possible, researchers need to access data. This sort of understanding is necessary for new standards-setting. There is an information asymmetry in which data is collected about users and non-users of platforms and algorithmic systems. But at the same, there is no access for researchers to understand the use of the collected d users and non-users of platforms and algorithmic systems is collected data about[89]. Therefore, an audit ecosystem needs leveled access by different auditors who have to publish and share their audit results. Such leveled access can reduce information asymmetry and, ultimately, power asymmetry.
4.5 Challenges Of Third-Party Audits
Recognizing the difficulties inherent in civil society and researcher audits, it is crucial to acknowledge that algorithm auditing represents only a single facet within the broader algorithm accountability framework. To establish genuine algorithm accountability, it is imperative to consider that algorithm auditing operates with various other factors, forming an intricate ecosystem [2]. This is why we demand the AI audit ecosystem to be open and interfere with other accountability processes. An essential component for that is actual enforcement, that is, post-audit actions such as Raji and Buolamwini [68]’s demands for harm reduction. This in turn requires European institutions that implement proper enforcement.
At the same time, one could say that p-hacking [8] allows third parties to select the metrics that indicate harm. However, this is not the same as the first-party audit case, as other researchers could quickly verify the harm. Civil society and social organizations would have no interest in making a bad name for themselves.
Furthermore, the AI Now Institute argues that third-party audits shift responsibility to researchers and affected communities. However, we cannot rely on these groups to have the resources to audit AI technologies [39] effectively. We see it as one measure that can work alongside other measures, such as audits through regulators, internal audits, transparency, and the aforementioned accountability processes.
While solving these problems presents considerable challenges, it is worth noting that the absence of civil society oversight would yield even more unfavorable outcomes. Adequate funding is essential to alleviate the burdens on civil society and research. Without such audits, achieving comprehensive accountability remains inadequate.
5 Addressing the Regulatory Gap: An EU AI Audit Ecosystem
Third-party audits by research and civil society are essential for an AI audit ecosystem. (1) Third parties can provide oversight as watchdogs against audit washing and false assurances. (2) Third-party audits can question standards, which could prevent the decoupling of audits, as in financial audits. (3) Third-party audits are adaptive to new technologies. In the context of (2) and (3), conducting audits through third-party entities gains further significance, as they can scrutinize evolving standards and raise critical inquiries [47]. (4) Third parties promote sensitivity to the sociotechnical system and (5) affected communities. (6) Last but not least, third-party audits and data access challenge information and power asymmetries of the platform economy. At the same time, we have shown that the DSA and AIA refer differently to third-party audits and permit different data access for third parties. We will assess these approaches to third parties and propose policy recommendations.
5.1 Evaluation of the Audit Status Quo in the DSA and the AIA
The DSA and the draft delegated act do not specify any additional requirements for the audit firm. [^77] The wording of Art. 37 DSA (“the organisations carrying out the audits”, Art. 37 (2) DSA) does not provide any further details on this. Thus, Art. 4 (1) Delegated Regulation on Independent Audits refers to Art. 37 DSA, according to which the provider must check whether the organisation to be selected meets the requirements of paragraph 3 before selecting the audit organisation. Furthermore, the Commission defines the term audit organisation means an individual organisation, a consortium or other combination of organisations, including any sub-contractors, that the audited provider has contracted to perform an independent audit in accordance with Article 37 DSA. This should be viewed critically for the two reasons: Firstly, it does not specify the requirements to be met by each criterion. For example, it does not explain in sufficient detail what level of “proven expertise” or supervisory steps are required to ensure compliance. In addition, the requirement of independence gives rise to discussion. [^78] For example, Spindler [81], who raises concerns, for example, that this does not clarify constellations in which an affiliate of the auditing company advises the provider Secondly, it is unclear whether audit firms need to be reviewed by bodies other than the VLOPs and VLOEs to ensure their proper functioning. However, as a regulatory instrument can only be as effective as the institutions that enforce it, the final version of the delegated regulation needs to set out more specific requirements. On the one hand, it should give priority to independent checks on the competent selection of auditors by setting out basic independence requirements. On the other hand, the regulation should provide for additional (official) supervision.
As noted above, the EU Commission’s AIA proposal relies primarily on internal conformity assessment as a means of assurance. The fact that the focus of the AIA is mainly on self-regulation is questionablebecause there is a lot at stake: high-risk AI systems covered by the AIA can violate numerous legal interests, e. g. people’s health and lives at risk, but also damage property or the environment or violate legal interests such as the protection of privacy. Moreover, implementing the risk-based approach of the AIA requires taking preventive measures that should not solely rely on the assessment of the AI provider, especially for high-risk systems [42].
Given the risks, whether a first-party internal audit can provide the necessary oversight of such potentially harmful high-risk AI systems is questionable. Despite regulatory requirements, enforcing internal conformity assessments is likely to be less stringent than it appears at first glance. In addition, conformity assessments that the provider voluntarily has carried out by third parties are likely to be rare in practice. Indeed, providers of high-risk systems are usually in the best position to understand and assess the conformity of the AI systems they develop. However, it is doubtful that first parties will seek to find all potential harms and thus find out their system does not comply with the AIA’s legal standards. We have laid out the issues of internal audits in the previous section.
For the above reasons, the requirements for audit firms in both the DSA and the AIA must be amended. The explicit inclusion of third parties would be beneficial under both acts. Under the current rules, there is a risk that providers will only engage second-party auditors to conduct mandatory annual assessments. This focus can not create an open, dynamic, and self-organizing ecosystem including various stakeholders. We have established the value of third-party audits for an AI audit ecosystem and shown that AI regulation in the EU must include third-party audits and their data access to account for appropriate oversight apart from compliance by first-party audits.
5.2 Advantages of Broader Data Access for (Vetted) Researchers
Purely relying on process access for third-party audits, as exemplified by the AMS algorithm in section 4.1.3, presents difficulties due to the inherent opacity of the system’s workings. This is why there is a growing call for a graduated system of access levels among various groups [44]. Consequently, we advocate for a minimum requirement to enable ’black-box’ access, preferably extended to non-vetted researchers without the need for formal accreditation. Reddit, Facebook, and Twitter opted to discontinue their APIs, which had been the authorized avenue for third parties to access and retrieve information regarding user activities on the platform. This situation exposes researchers to legal liabilities due to potential breaches of, for example, Facebook’s Terms of Service, as previously elucidated in the context of AlgorithmWatch’s data donation case [50].
The research efforts undertaken in this context should be safeguarded by robust safe-harbor provisions that grant legal immunity to these researchers thereby shielding them from potential legal charges linked to hacking or unauthorized access, as the AI Now Institute demands [39]. This framework would offer protection to those conducting audits that potentially expose harm, thereby incentivizing such assessments. More widespread access to research data can, to some extent, reduce this power imbalance through ”audit capture” [46]. Broader access to data should prevent such scenarios of audit capture, seen when Facebook halted AlgorithmWatch’s research on Instagram; this emphasizes the role of third-party audits and other measures to reduce market power.
Consequently, one fundamental proposition is the establishment of a right to scrape, although an even more effective approach could involve mandatory APIs. However, implementing the latter may pose challenges for non-vetted researchers, as it might inadvertently allow them to utilize the product without incurring costs. For vetted researchers, broader access should be granted, enabling performance assessments akin to the approach taken in the gender shades case and possibly allowing the measurement of facial recognition performance. We acknowledge that the DSA includes such access for vetted researchers.
We have seen that NGOs have can ensure social accountability and oversight trough domain knowledge and inclusion of affected people. At the same time, journalists have already shown in the past how important they are as an accountability component in financial audits and also in algorithmic audits. It is for this reason that we believe that the extension of vetted researchers to include NGOs and journalists makes sense. This extension could be done in the delegated acts by describing vetted researchers in a broad manner such that NGOs and journalists are included.
As has been shown, the AIA does not contain any provisions comparable to the DSA that facilitate regulated access by researchers to data held by AI providers. This is surprising given that, according to the rationale for its adoption, the AIA should in principle facilitate this. We recommend that this position be reconsidered and added to the final document.
5.3 Policy Recommendations
Explicit Inclusion of Third Parties
Future AI regulations should explicitly include provisions for third-party audits and third-party access by researchers, civil society, and regulators. This ensures a more comprehensive and independent assessment of AI products and platform and moves towards establishing an EU AI audit ecosystem.
Inclusion of NGOs and Investigative Journalists
In the delegated acts of the DSA, the inclusion NGOs and investigative journalists as vetted researchers who can perform third-party audits is necessary. Their involvement can ensure social accountability through domain knowledge and special access to accountability by including affected communities. At the same time, journalists have already shown in the past how important they are as an accountability component in financial audits and also in algorithmic audits.
Third-Party Access in AIA
Incorporate provisions for third-party access in the AIA, particularly for ”high-risk” AI systems. These provisions ensure that critical systems undergo rigorous scrutiny and that hard standards can be questioned in the light of new technologies.
Incentivize Third-Party Audits
Create incentives, such as funding or tax benefits, to encourage organizations to undergo third-party audits voluntarily. These incentives can be funded through companies managed by the government or government funding. This promotes an enabling environment for an AI audit ecosystem and thus the promotion of algorithmic conformity assessment procedures within the framework of all regulations. The government could choose the researchers eligible for funding in the case of the DSA (should be thought of in the delegated acts) and through a multistakeholder board in the AIA which includes state-run ADMs.
Publish Audit Results
Nondisclosures of audit results should be an exception if sensitive information or trade secrets are undeniably part of the audit results. Otherwise, mandate the publication of anonymized audit results, whether conducted by first, second, or third parties. This can inform other auditors and promote further knowledge sharing.
Harm Incident Reporting and Detection
Implement measures to detect and report AI-related harms. Establish a framework for reporting incidents and ensure post-audit actions and enforcement are taken.
Whistleblower Protections in AI Companies
Whistleblower protections for employees who expose violations of AI standards, legal and regulatory requirements, civil rights, and human rights should be strengthened. Whistleblowers should be shielded from retaliation. The EU Whistleblower Protection Directive [25] is a step in the right direction and will be welcomed. It will be compelling to see how AIA and the EU Whistleblower Protection Directive interact.
Safe Harbor Provisions for Auditing Entities
Implement safe harbor provisions specifically for researchers, NGOs, and investigative journalists who conduct AI audits. These entities should be granted legal protection from retaliatory actions when conducting audits in good faith. Safe harbor provisions should apply when audits are conducted according to established standards and guidelines and findings are reported accurately.
Legalization of Web Scraping for AI Auditing
Recognize the importance of web scraping as a valuable tool for researchers, NGOs, and investigative journalists engaged in AI auditing activities. Establish clear legal frameworks that permit responsible web scraping when it is carried out for legitimate research and auditing purposes. These frameworks should include safeguards to protect against misuse and ensure the ethical and responsible use of web scraping techniques.
6 Conclusion
We defined audit types – focusing on third-party audits – and analyzed to which extent these audit types and data access are included in the EU regulatory framework.
According to our analysis, the DSA will establish a novel institutional system that involves third-party audits. At the same time, the DSA provides data access for vetted researchers. However, the requirements for independence in the DSA still need to be specified. The European legislator’s AIA proposal relies primarily on internal conformity assessment. At present, access for researchers is not explicitly regulated.
Based on this analysis, we define an AI audit ecosystem that accounts for compliance and oversight and demonstrate the existence of a regulatory gap. We emphasize that third-party audits by research and civil society must be part of that ecosystem. We demand that the AIA include data and model access for certain AI products.
We hope to contribute to the emerging field of audits as a means of accountability. In doing so, we aim to sharpen the focus on affected communities and incorporate past lessons from platform research and financial audits. We recommend establishing an AI audit ecosystem to ensure compliance and oversight. We concur with Raji et al. [70] demands for an ”ecosystem in which third parties can not only survive in their role, but thrive in directly confronting, verifying, and subjecting to scrutiny the performance claims made by corporations, while adequately addressing complaints of harm brought forth by the impacted population.” Our policy proposals point to democratizing accountability, and we hope future AI regulations and the EU directives mentioned here will address them.
7 Compliance with Ethical Standards
7.1 Conflict of Interest
On behalf of all authors, the corresponding author states that there is no conflict of interest.